A car starting from rest accelerates at the rate of f through distance S then continues at constant speed for tome t and decelerates at the rate f/2 to come to rest .if total distance travelled is 15s then S
1
Expert's answer
2019-08-12T10:02:32-0400

The constant speed is


"v^2=2fs"

"v=\sqrt{2fs}"

Distance for the second period is


"D=vt=\sqrt{2fs}t"

For the third period:


"v^2=2(0.5f)d""2fs=fd"

"d=2s"

Total distance travelled is 

"15s=s+\sqrt{2fs}t+2s"

"\sqrt{2fs}t=12s"

Thus,


"s=\frac{ft^2}{72}"